Does there exist a plane targent to x^2−2y^2 +2z^2 = 8 and which passes through 2x+3y+2z = 8, x−y+2z = 5? Justify your answer.
1
Expert's answer
2020-02-17T11:25:28-0500

"\nabla(F)=(F_x,F_y.F_z)=(2x,-4y,4z)"

Let's find the intersection of two planes, which should be a line with a directional vector equal to the cross product of normal vectors of these planes


"v=(2,3,2)\times(1,-1,2)=(8,-2,-5)"

Since the plane we are looking for, passes through that line, it's normal vector of the plane should be parallel to the line's directional vector, therefore

"2x=8;\quad
-4y=-2;\quad
4z=-5"

"x=4;\quad y=\frac{1}{2};\quad z=-\frac{-5}{4}"

The equation of the tangent plane

"8(x-4)-2(y-\frac{1}{2})-5(z+\frac{5}{4})=0"